1/3+(-5/6) + (-1/2) =?
{-27+14}=?
-1/2-(5/8)=?
-3.4-(7.6)=?

Answer by checkley75(3666)

1/3+(-5/6)+(-1/2)=1/3-5/6-1/2
now we need to find a common denominator for 3,6 & 2. 6 is the LCD.
(1*2-5-3*1)/6=(2-5-3)/6=-6/6=-1
-----------------------------------------------
-27+14=-13
-------------------------------------------------
-1/2-5/8 the common denominator hjere is 8
(-1*4-5)/8=(-4-5)/8=-9/8=-1 1/8
-------------------------------------------------
-3.4-7.6=-11